OpenStudy (amirahlovespink):

Artie needs to save $450 to buy the new bicycle that he wants. He currently has $300. What percent of the cost of a new bicycle has he saved? Claire decides that to solve this problem by setting up a proportion. What should her proportion look like?

OpenStudy (orange_):

Proportions are just equal fractions. Can you come up with a guess? I can help you. :D

OpenStudy (amirahlovespink):

A. 300/100=x/450 B.X/100=300/450 C.X/100=450/300 D. 450/100=X/300

OpenStudy (amirahlovespink):

EH.... i sorta dont get it

OpenStudy (orange_):

Sorry, I don't either.

OpenStudy (orange_):

Look up proportion